# Scientific Machine Learning of Chaotic Systems Discovers Governing Equations for Neural Populations

### v2

@@ IV.5 Experimental data acquisition and preprocessing
As acquisition and preprocessing of each of the three experimental datasets have been described in detail in their respective publications [63, 65, 67], we only briefly describe the datasets here. The rat MEC dataset [64] was acquired from six male Long-Evans rats, with electrode recordings acquired during a long period of roaming in a circular track. The rat prefrontal cortex dataset [66] was acquired from four male Long-Evans rats during 10 trials in which they successfully learned a behavioral task, including the rest periods before and after the task recordings. The human orbitofrontal cortex dataset [68] was acquired during electrocorticography recordings of 10 neurosurgical patients undergoing evaluation for epilepsy. The patients were asked to play a decision-making game, and the recordings used for this data are the inter-trial rest portions of the ECoG data. For all datasets, the synchrony between electrodes within the region was computed as the phase-locking value (PLV) between the angle of Hilbert-transformed signals ($\phi$) from each electrode [87, 88]. Following prior work [48], spike data from the dataset, which recorded spike times rather than LFPs, were converted to an LFP proxy by windowed spike counts, with the PLV calculation performed on these instead. PLV is given by the formula
@@ VI Code and Data Availability
All code and simulated data to reproduce the results and plots in this manuscript, along with an additional abbreviated demo version for interested readers to rapidly prototype with, can be found in the GitHub repository associated with this publication at https://github.com/Neuroblox/pem-ude. The intracranial electrode recording datasets can all be accessed in repositories on CRCNS.org, with the specific conditions of access to each one listed on their individual repository pages [64, 66, 68].

### v3

@@ IV.5 Experimental data acquisition and preprocessing
As acquisition and preprocessing of each of the three experimental datasets have been described in detail in their respective publications [hernandez-perez2020, peyrache2009, saez2018], we only briefly describe the datasets here. The rat MEC dataset [hernandez-perez2020a] was acquired from six male Long-Evans rats, with electrode recordings acquired during a long period of roaming in a circular track. The rat prefrontal cortex dataset [peyrache2018] was acquired from four male Long-Evans rats during 10 trials in which they successfully learned a behavioral task, including the rest periods before and after the task recordings. The human orbitofrontal cortex dataset [saez2018a] was acquired during electrocorticography recordings of 10 neurosurgical patients undergoing evaluation for epilepsy. The patients were asked to play a decision-making game, and the recordings used for this data are the inter-trial rest portions of the ECoG data. For all datasets, the synchrony between electrodes within the region was computed as the phase-locking value (PLV) between the angle of Hilbert-transformed signals ($\phi$) from each electrode [kuramoto1984, mormann2000]. Following prior work [pathak2024], spike data from the dataset, which recorded spike times rather than LFPs, were converted to an LFP proxy by windowed spike counts, with the PLV calculation performed on these instead. PLV is given by the formula
@@ VI Code and Data Availability
All code and simulated data to reproduce the results and plots in this manuscript, along with an additional abbreviated demo version for interested readers to rapidly prototype with, can be found in the GitHub repository associated with this publication at https://github.com/Neuroblox/pem-ude. The intracranial electrode recording datasets can all be accessed in repositories on CRCNS.org, with the specific conditions of access to each one listed on their individual repository pages [hernandez-perez2020a, peyrache2018, saez2018a].
@@ VII Competing Interests
VIII Supplemental Information
VIII.1 Supplemental Contents
Four supplemental figures: Fig. 6 for a complete description of the method presented here, Fig. 7 for the supplemental parameter profiles of the Rössler system under PEM, Fig. 8 for the 2D parameter landscapes of the Rössler system under PEM and Fig. 9 for the influence of PEM on the divergence of shadow trajectories.
VIII.2 Refining Symbolic Regression for NGNMM discovery
We provide additional notes on the fitting of the PEM-UDE system for the NGNMM at different sparsities, as well as comments on the terms learned by STLSQ in the symbolic form of the system. First, we note that the PEM-UDE approach is able to learn well enough to generalize to sparsities not originally in the training data. As an example, in Fig. 10 we replot Fig. 4e, but using a network trained only on $p_{c}=0.3$ to $p_{c}=0.9$ data. We note that even though this is limited to intermediate degrees of sparsity, the PEM-UDE system still adequately recovers the dynamics down to even 5% connectivity ($p_{c}=0.05$), albeit with a slight frequency shift. While we do not exploit this feature to its full potential in this paper, we note this as a promising future direction for further applications of the PEM-UDE approach, which synergizes with recent approaches that demonstrate neural ODEs are able to generalize beyond bifurcations unseen in the training data [van_tegelen_neural_2025].
The primary advantage of fitting a symbolic form rather than using the UDE black-box dynamics is the ability to analyze the learned terms. This is particularly true for comparing the new form of the NGNMM with other analytical approaches that have attempted to address mean-field activity of incompletely connected networks of neurons [divolo2018, divolo2022] or networks driven by distributed noisy processes [goldobin2019]. As an illustration of this approach, we further consider the dynamics learned from the $p_{c}=0.5$ network shown in Fig. 11. Training of the UDE was carried out in the following way: 1) The UDE (eqs 22) is trained using a fixed $p_{c}=0.5$, on the simulated firing rate $r$ and average voltage $v$ with $K=0.3$ using the Adam optimizer for 40,000 iterations. We then add shorter training steps during which we slowly reduce $K$ to $0.1$. We found that below $K=0.1$, the optimization does not improve further. The reason for reducing $K$ is to mitigate any residual influence of the PEM term on the weights of the neural network and to ensure that UDE reproduces the experimental data without the PEM term. Applying symbolic regression to the resulting UDE, we find similar terms to those discovered for $\phi_{1}$ - $\phi_{6}$ in Eq. 23, that is, a dependence of firing rate upon synaptic activity ($\phi_{1}$) and voltage-dependent synaptic activity ($\phi_{3}$), along with an introduction of a linear dependence on $r$ modulated by voltage and synaptic dynamics ($\phi_{4}$ - $\phi_{6}$). The approach presented here differs most from other derivations in the discovery of a novel quadratic dependence of voltage upon synaptic dynamics ($\phi_{7}$). While this is not present in the simpler network structure, it is an emergent feature of supralinear dependence on dendritic integration often seen in biological networks (e.g., NMDA receptor integrated signaling [poirazi_illuminating_2020]) and used in more complicated neuronal models that focus specifically on dendritic dynamics [abbott_realistic_1991]. Future models should incorporate these dynamics at the population level as well to provide a better ground-truth training data for this approach.
Figure 6: Overview of the entire pipeline. This is a more explicitly descriptive version Fig. 1.
Figure 7: Loss landscapes of the Rössler system parameters not fit by the PEM-UDE process as the error correction is applied to the second state. We note that for both parameters $b=0.22$ (a) and $c=14$, the addition of the correction term to the second state (where neither appears) tends to smooth their loss landscapes but create a plateau around the true value of the parameter as the correction term is increased. This means that the PEM-UDE fitting method does not fundamentally alter the landscape in a way that would lead to incorrect dynamics, but that the accuracy of the final fit will depend on how well the parameters not fit by the PEM-UDE are estimated.
Figure 8: Loss landscapes for both the fit ($a$) and unfit ($b$, $c$) parameters when the PEM method is applied to the second state (which contains only $a$). Notice that with increasing smoothing $K$, the global minimum becomes more tractable to find via gradient descent, but there is a ridge near the minimum that will require either knowledge or a good estimate from data of the parameters not being fit to ensure accuracy in the final fit.
Figure 9: Sufficiently large hyperparameter $K$ for the PEM removes sensitive dependence on initial parameter values. a. The trajectories of the Rössler system diverge for two very close parameter values ($a=0.2$ vs $a=0.200000001$), but this divergence is mitigated by sufficiently large $K$ in the PEM correction term. b. This correction provides the same aid to fitting for any reasonable deviation from the true parameter (shown for deviations in the range [1e-12, 1e-1].
Figure 10: The PEM-UDE learned dynamics, even when constrained during training to samples representing only $p_{c}=0.3$ - $p_{c}=0.9$, still learn dynamics that generalize down to greater sparsities (e.g., $p_{c}=0.05$).
Figure 11: STLSQ-derived symbolic terms presented in Eq. 23 accurately recover dynamics at $p_{c}=0.5$. See Supplemental Information B for a further discussion of how these terms map onto other derivations and biological networks.